I AM HAVING PROBLEMS SOLVING THIS PROBLEM AND NEED SOME HELP
SOLVE BY COMPLETING THE SQUARE
X^2=5X+2
THANK YOU

Answer by stanbon(75887)

SOLVE BY COMPLETING THE SQUARE
X^2=5X+2
------------
x^2-5x+? = 2+?
Complete the square to get:
x^2-5x+[(1/2)*(-5)]^2 = 2 + [(1/2)*(-5)]^2
Simplify to get:
x^2 - 5x + (25/4) = 2 + (25/4)
Factor to get:
(x-(5/2))^2 = 33/4
Take the square root of both sides to get:
(x-(5/2)) = [sqrt(33)]/2 or (x-(5/2)) = [-sqrt(33)]/2
x = (5/2)+(sqrt(33)/2 or x = (5/2) - [sqrt(33)]/2
===================
Cheers,
Stan H.